1) Together, Max and Raoul sold 137 tickets for a benefit baseball game. If Max sold 10 fewer than twice as many as Raoul, how many tickets did each boy sell?
2) Mary and Anthony are driving 1780 miles from New York to Denver. If their car gets 25 mil/gal and gasoline averages about $1.03/gal, how much money will they pay for gas on their trip?

Answer by jorel1380(3719)

n+(2n-10)=137
3n-10=137
3n=147
n=49
Raoul sold 49 tickets; Max sold 88 tickets..
-----------------------
1780/25x1.03=71.2x1.03=$73.336 for gas..